# Scaling Laws for Emulation of Stellar Spectra

### v1

@@ 1 Introduction
Several approaches have been developed to address these needs, including quadratic models such as The Cannon (Ness et al., 2015), a polynomial-based ridge regression; neural network-based models such as The Payne (Ting et al., 2019); and TransformerPayne (Różański et al., 2024).
@@
The concept of scaling laws has also begun to be explored in astronomical research, including applications to stellar light curves (Pan et al., 2024) and galaxy images (Walmsley et al., 2024; Smith et al., 2024). Demonstrating the validity of scaling laws represents a critical step toward developing deep learning foundational models for astronomy. These foundational models are characterized by their broad generalization and adaptability at a given model size, similar to how large language models, while primarily trained on English text, can generalize to respond in other languages with relatively limited examples, demonstrating effective domain transfer capabilities.
@@ 2 Methods
The goal of this study is to understand how emulator quality evolves as we scale to larger training datasets, larger neural networks, and longer training durations. We formulate the emulation problem following our previous work on the TransformerPayne model (Różański et al., 2024).
In these models, we construct a function $f_{\theta}(\lambda,\mathbf{p})$ which approximates exact normalized flux, where $\lambda$ represents the wavelength and $\mathbf{p}$ denotes a collection of stellar properties, including stellar atmosphere parameters, such as the effective temperature ($T_{\text{eff}}$) or the logarithm of surface gravity ($\log g$), and elemental abundances. In this work, the parameter vector $\mathbf{p}$ consists of one hundred parameters, including effective temperature, surface gravity, and elemental abundances spanning from helium to einsteinium. This function is implemented as a neural network parameterized by free parameters $\theta$, which are optimized to make $f_{\theta}(\lambda,\mathbf{p})$ an accurate approximation of spectra across the domain of interest.
@@
Figure 1: Architecture diagram of the TransformerPayne variant implemented for our scaling experiments. The model processes two inputs: wavelength (lower left) and spectrum parameters (lower right). These inputs are transformed through Sinusoidal Embedding and MLP Embedding respectively, then normalized via RMSNorm. The embedded representations flow through N sequential Transformer blocks (two are explicitly shown here for clarity), each containing Feed Forward networks and Multi-Head Attention mechanisms with residual connections. Each block processes outputs from the previous block and maintains the information flow through normalization layers. The final representation is processed by an MLP Head to predict the normalized flux. This architecture incorporates the Pre-LN placement scheme for residual connections, which facilitates stable training dynamics when scaling to larger model configurations.
@@ 2.1 MLP and Sinusoidal Embedding
where $\omega_{i}$ follow a geometric progression covering a manually adjusted range of circular frequencies, from $\omega_{\text{min}}$ to $\omega_{\text{max}}$, and index $i$ from 1 to $d$. We choose frequencies that span from those corresponding to broad spectral features down to fractions of the narrowest line profiles, ensuring comprehensive coverage of spectral characteristics at multiple scales.
@@ 2.2 Transformer Block
$\mathbf{r}_{i}$ $=\;\mathbf{x}\;+\;\mathbf{MHA}\!\Bigl{(}\text{RMSNorm}(\mathbf{x}),\,\text{RMSNorm}(\mathbf{p_{\text{emb}}})\Bigr{)},$ (11)
$\mathbf{r}_{i}$ $=\;\mathbf{r}_{i}\;+\;\text{FF}\!\bigl{(}\text{RMSNorm}(\mathbf{r}_{i})\bigr{)},$
@@ 2.3 Dimensions of Scaling Laws
First, we define the size of the training data as the number of distinct spectra, that is, spectra computed using different sets of atmospheric parameters. These training samples are drawn from a uniform distribution spanning the entire parameter space of interest, see Sec. 2.4 for exact ranges. This approach ensures comprehensive coverage of the stellar parameter domain being modeled.
@@ 2.4 Spectral Grid, Training and Metrics
The experiments in this work utilize one of the synthetic datasets of 100,000 normalized stellar spectra introduced in Różański et al. (2024). These spectra were generated using plane-parallel Local Thermodynamic Equilibrium (LTE) atmospheric models (Lester and Neilson, 2008; Kurucz, 1979, 2005). The wavelength range spans from 4000 to 5000 Å at a resolution of $R=100,000$, sampled at 22,315 wavelengths spaced equidistantly. The grid varies the effective temperature, $T_{\text{eff}}$, from 4000 K to 6000 K and the surface gravity from 4 to 5. The microturbulence velocity remains fixed at $\xi=0$ km/s. Helium abundances vary between none and twice the solar value (assuming 0.0782 in the Sun), while the abundances of other elements (atomic numbers $Z=3$ to $Z=99$) are sampled uniformly between $-2$ and 1 dex relative to the solar abundance. It is important to note that these spectra are normalized using the theoretical continuum, resulting in flux values predominantly near 1 (with absorption features appearing as downward deviations below 1). Consequently, all MSE values reported in this study are with respect to this normalized scale, making them dimensionless quantities that directly represent fractional deviations in the normalized flux.
@@
The WSD schedule is particularly useful for scaling law experiments, as it allows the warm-up phase to be fixed while deferring the cool-down phase until later in training. This flexibility significantly reduces the time required to test different training durations without incurring substantial computational overhead. In our initial tests, we verified that WSD performs at least as well as the linear warm-up followed by cosine decay schedule used in the TransformerPayne paper (Różański et al., 2024), making it a suitable choice for our study. Additionally to global learning rate $\eta$, we also tuned an embedding’s learning rate scaling factor $\eta_{\text{E}}$.
@@ 2.6 Neural Scaling Laws
Following the convention established by Kaplan et al. (2020), we consider three primary scaling regimes, each corresponding to a different limiting factor:
@@ 3.4.4 Optimal Resource Allocation
While the right panels of Fig. 6 show how MSE scales with dataset size and model size independently, what is often more practically valuable is understanding how to optimally balance resources when compute is the primary constraint. This is particularly important because the largest models are not always optimal, at limited compute budgets, smaller models trained for more steps often outperform larger models trained insufficiently. Fig. 7 addresses precisely this practical question.
@@ 4.4 Caveats and Future Work
We focus on MSE as our primary metric because emulation errors manifest as systematic noise in the spectral fitting process, which, to a first approximation, can be translated into statistical errors on stellar parameters using Fisher matrix calculations. This direct connection between spectral emulation quality and parameter inference precision makes MSE a robust and theoretically justified optimization target (Ting et al., 2017; Sandford et al., 2020; Różański et al., 2024).

### v2

@@ 1 Introduction
Several approaches have been developed to address these needs, including quadratic models such as The Cannon (Ness et al., 2015), a polynomial-based ridge regression; neural network-based models such as The Payne (Ting et al., 2019); and TransformerPayne (Różański et al., 2025).
@@
The concept of scaling laws has also begun to be explored in astronomical research, including applications to stellar light curves (Pan et al., 2024) and galaxy images (Walmsley et al., 2024; Smith et al., 2024). Demonstrating the validity of scaling laws represents a critical step toward developing deep learning foundational models for astronomy. Scale plays a crucial role, as the generalizability of such foundational models primarily emerges with increasing scale. Analogously, large language models trained predominantly on English text exhibit improved ability to adapt to new languages or specialized domains when their scale is increased, highlighting scale-dependent improvements in domain transfer capabilities.
@@ 2 Methods
The goal of this study is to understand how emulator quality evolves as we scale to larger training datasets, larger neural networks, and longer training durations. We formulate the emulation problem following our previous work on the TransformerPayne model (Różański et al., 2025).
In these models, we construct a function $f_{\theta}(\lambda,\mathbf{p})$ which approximates exact normalized flux, where $\lambda$ represents the wavelength and $\mathbf{p}$ denotes a collection of stellar properties. In this work, the parameter vector $\mathbf{p}$ consists of one hundred parameters, including effective temperature, surface gravity, and elemental abundances spanning from helium to einsteinium. This function is implemented as a neural network parameterized by free parameters $\theta$, which are optimized to make $f_{\theta}(\lambda,\mathbf{p})$ an accurate approximation of spectra across the domain of interest.
@@
Figure 1: Architecture diagram of the TransformerPayne variant implemented for our scaling experiments. The model processes two inputs: wavelength (lower left) and spectrum parameters (lower right). These inputs are transformed through Sinusoidal Embedding and MLP Embedding respectively, then normalized via RMSNorm. The embedded representations flow through N sequential Transformer blocks, each containing Feed Forward networks and Multi-Head Attention mechanisms with residual connections. Two Transformer blocks are shown for clarity; the ellipsis symbol and disconnections near $\mathbf{r_{1}}$ indicate that the remaining $N-2$ blocks are omitted from the figure. The circled ”+” symbols represent element-wise summation in residual connections. Each block processes outputs from the previous block and maintains the information flow through normalization layers. The final representation is processed by an MLP Head to predict the normalized flux. This architecture incorporates the Pre-LN placement scheme for residual connections, which facilitates stable training dynamics when scaling to larger model configurations.
@@ 2.1 MLP and Sinusoidal Embedding
where $\omega_{i}$ follow a geometric progression covering a manually adjusted range of circular frequencies, from $\omega_{\text{min}}$ to $\omega_{\text{max}}$, and index $i$ from 1 to $d$. We select these frequencies to cover all relevant scales, from continuum-level variations across the full 4000–5000 Å interval down to features smaller than the narrowest lines. Concretely, working in $\log_{10}\lambda$ ( Å), we take periods, $P$, geometrically spaced from 10 to $10^{-6}$ and convert them to angular frequencies via $\omega_{i}=2\pi/\text{P}_{i}.$ The longest period tracks more than the entire spectral window, while the shortest resolves scales of order 0.05 Å and smaller - about a resolution of our spectral grid. This configuration performs well in practice, as confirmed by initial experiments, although fine-tuning the frequency range or spacing scheme could further improve emulation accuracy.
@@
To summarize, the embedding choices reflect domain-specific inductive biases tailored to the characteristics of the stellar spectra. Wavelengths are mapped using a fixed sinusoidal embedding to provide high-frequency basis functions that support the representation of sharp spectral features without increasing model complexity. By contrast, stellar parameters influence the flux more smoothly, so a lightweight two-layer MLP embedding is sufficient to capture their effects. While effective in practice, these choices are not necessarily optimal; more expressive embeddings may offer further gains and are a subject for future exploration.
@@ 2.2 Transformer Block
$\mathbf{\hat{r}}_{i}$ $=\;\mathbf{x}\;+\;\mathbf{MHA}\!\Bigl{(}\text{RMSNorm}(\mathbf{x}),\,\text{RMSNorm}(\mathbf{p_{\text{emb}}})\Bigr{)},$ (11)
$\mathbf{r}_{i}$ $=\;\mathbf{\hat{r}}_{i}\;+\;\text{FF}\!\bigl{(}\text{RMSNorm}(\mathbf{r}_{i})\bigr{)},$
@@ 2.3 Dimensions of Scaling Laws
First, we define the size of the training data as the number of distinct spectra, that is, spectra computed using different sets of atmospheric parameters. These training samples are drawn from a uniform distribution spanning the entire parameter space of interest, see Sec. 2.4 for exact ranges. A uniformly sampled grid is straightforward to construct and is often sufficient for training; however, in regions of the parameter space that exhibit steep gradients, a locally denser or otherwise adaptive sampling strategy may be preferable. Such schemes may substantially improve sample efficiency and are left for future investigation.
@@ 2.4 Spectral Grid, Training and Metrics
The experiments in this work utilize one of the synthetic datasets of 100,000 normalized stellar spectra introduced in Różański et al. (2025). These spectra were generated using plane-parallel Local Thermodynamic Equilibrium (LTE) atmospheric models (Lester and Neilson, 2008; Kurucz, 1979, 2005). The wavelength range spans from 4000 to 5000 Å at a resolution of $R=100,000$, sampled at 22,315 wavelengths spaced equidistantly. The grid varies the effective temperature, $T_{\text{eff}}$, from 4000 K to 6000 K and the surface gravity from 4 to 5. The microturbulence velocity remains fixed at $\xi=0$ km/s. Helium abundances vary between none and twice the solar value (assuming 0.0782 in the Sun), while the abundances of other elements (atomic numbers $Z=3$ to $Z=99$) are sampled uniformly between $-2$ and 1 dex relative to the solar abundance. Although Różański et al. (2025) showed that only 38 elemental abundances are well constrained on this spectral grid, we retain the full training dataset to ensure direct comparability with the results presented in that work. Adding parameters with no spectral effect does not compromise accuracy, as the model internally discards them through learned representations. It is important to note that these spectra are normalized using the theoretical continuum, resulting in flux values predominantly near 1 (with absorption features appearing as downward deviations below 1). Consequently, all MSE values reported in this study are with respect to this normalized scale, making them dimensionless quantities that directly represent fractional deviations in the normalized flux.
@@
The WSD schedule is particularly useful for scaling law experiments, as it allows the warm-up phase to be fixed while deferring the cool-down phase until later in training. This flexibility significantly reduces the time required to test different training durations without incurring substantial computational overhead. In our initial tests, we verified that WSD performs at least as well as the linear warm-up followed by cosine decay schedule used in the TransformerPayne paper (Różański et al., 2025), making it a suitable choice for our study. Additionally to global learning rate $\eta$, we also tuned an embedding’s learning rate scaling factor $\eta_{\text{E}}$.
@@ 2.6 Neural Scaling Laws
Following the convention established by Kaplan et al. (2020), we consider three primary scaling regimes, each representing a scenario in which emulator performance is capped by a single scarce resource – model parameters, training data, or training compute:
@@ 3.4.4 Optimal Resource Allocation
While the right panels of Fig. 6 show how MSE scales with dataset size and model size independently, what is often more practically valuable is understanding how to optimally balance resources when compute is the primary constraint. This point is particularly important because the largest models are not always optimal. Under limited compute budgets, smaller models trained for more optimization steps often outperform larger models trained for fewer steps. Fig. 7 addresses precisely this practical question.
@@ 4.4 Caveats and Future Work
We focus on MSE as our primary metric because emulation errors manifest as systematic noise in the spectral fitting process, which, to a first approximation, can be translated into statistical errors on stellar parameters using Fisher matrix calculations. This direct connection between spectral emulation quality and parameter inference precision makes MSE a robust and theoretically justified optimization target (Ting et al., 2017; Sandford et al., 2020; Różański et al., 2025).